I came up with parallel on this problem. Please check my work.
Are the following lines parallel,perpendicular, or neither
L1 through (-4,-7) and (1,3)
L2 through (2,6) and (4,10)
Thank you for your assistance,
JRR

Answer by Nate(3500)

(-4,-7) and (1,3)
M = (3 + 7)/(1 + 4) = (10)/(5) = 2
(2,6) and (4,10)
M = (10 - 6)/(4 - 2) = (4/2) = 2
Great Job
You Are Right!